# Convolutional Neural Networks and Mixture of Experts for Intrusion Detection in 5G Networks and beyond

### v1

@@ Abstract
The advent of 6G/NextG networks comes along with a series of benefits, including extreme capacity, reliability, and efficiency. However, these networks may become vulnerable to new security threats. Therefore, 6G/NextG networks must be equipped with advanced Artificial Intelligence algorithms, in order to evade these attacks. Existing studies on the intrusion detection task rely on the train of shallow machine learning classifiers, including Logistic Regression, Decision Trees, and so on, yielding suboptimal performance. Others are based on deep neural networks consisting of static components, which are not conditional on the input. This limits their representation power and efficiency. To resolve these issues, we present the first study integrating Mixture of Experts (MoE) for identifying malicious traffic. Specifically, we use network traffic data and convert the 1D array of features into a 2D matrix. Next, we pass this matrix through convolutional neural network (CNN) layers followed by batch normalization and max pooling layers. After obtaining the representation vector via the CNN layers, a sparsely gated MoE layer is used. This layer consists of a set of experts (dense layers) and a router, where the router assigns weights to the output of each expert. Sparsity is achieved by choosing the most relevant experts of the total ones. Finally, we perform a series of ablation experiments to prove the effectiveness of our proposed model. Experiments are conducted on the 5G-NIDD dataset, a network intrusion detection dataset generated from a real 5G test network. Results show that our introduced approach reaches weighted F1-score up to 99.95% achieving comparable performance to existing approaches. Findings also show that our proposed model achieves multiple advantages over state-of-the-art approaches.
@@ I Introduction
Existing studies train shallow machine learning classifiers, obtaining in this manner suboptimal performance. Recently, existing studies have moved their interests in converting network traffic data or packet-level data into images and then training convolutional neural networks (CNNs), pretrained CNNs (AlexNet, VGG19, ResNet), RNNs (LSTMs, BiLSTMs), and after that employing fully-connected layers being responsible for classification. However, these approaches depend on dense layers being responsible for doing everything. Network parameters are fixed during training, while inference is performed in a static manner, demanding in this way computational resources and increasing both training and inference time. On the contrary, literature review suggests that models conditioned on the input [5] enjoy a series of benefits, including efficiency, representation power, adaptiveness, compatibility. This is attributable to the fact that these models selectively activate model components. Mixture of Experts belong to the category of the input-conditional computation models. Specifically, MoE were originally proposed in [6] and since then they have found applications in a series of domains [7, 8, 9].
To address the aforementioned limitations, we present the first study integrating MoE layers into a deep neural network in the task of intrusion detection in 5G networks. Specifically, we use network traffic data consisting of a feature set, which corresponds to a 1D array. We reshape this array into a matrix, which is then given as input to CNN, max-pooling, and batch normalization layers. After this, we use a sparsely-gated MoE layer [9], which applies different subsets of layers (experts) and activates only a selected subset of experts, i.e., the k most relevant ones, during each forward pass. Experiments are performed on the 5G-NIDD dataset [10], which contains 8 types of attacks. Results demonstrate that the proposed method achieves notable benefits over state-of-the-art approaches.
@@ III Dataset
Data of 5G-NIDD are extracted from two base stations. The authors in [10] have made available a csv file, which corresponds to the combined dataset, consisting of network flow data. 5G-NIDD dataset contains two attack categories, namely DoS/DDoS and Port Scan attacks. In terms of the DoS/DDoS attacks, the following types are included: ICPM flood, UDP flood, SYN flood, HTTP flood, and Slow rate DoS. Regarding the Port Scan attack, the following types are included: SYN Scan, TCP Connect Scan, and UDP Scan.
IV Methodology
Figure 1: Proposed Methodology
(a) CNN Architecture ($f_{\theta}$)
(b) CNN cell structure
Figure 2: Description of $f_{\theta}$
IV-A Preprocessing
IV-A1 Handling Missing Values
@@ IV-A2 Designing the feature set
TABLE I: Description of features
@@
IV-B Convolutional Neural Networks
@@
We design a convolutional neural network architecture to obtain a representation vector from input $X$. The CNN component of our methodology is illustrated in Fig. 2a. As one can observe, we place 4 CNN cell in a row with 16, 32, 64, and 128 filters respectively. Kernel size and padding are set equal to 3 and 1 respectively.
@@
The structure of the CNN cell is illustrated in Fig. 2b. Specifically, each CNN cell, except for the final one with 128 filters) consists of an 1D Convolutional, Batch normalization, ReLU activation, and Max-Pooling layer. The final CNN cell does not include a max pooling layer. In terms of the max pooling layer, we set both kernel size and stride equal to 2.
@@
IV-C Sparse Mixture of Experts
@@ V-B Experimental Setup
We use 128 (value of $n$) experts. We keep the 32 (value of $k$) most relevant experts. We set $\alpha$ of Eq. 10 equal to 0.1. We use a batch size of 1024. We divide the dataset into a train and test set (60-40%) in a stratified manner. We train the models for a maximum of 40 epochs. We use PyTorch [33] for performing our experiments. All experiments are performed on a NVIDIA A100 80GB PCIe GPU.
@@ V-C Evaluation Metrics
We also report weighted F1-score, since we have used an imbalanced dataset.
@@ V-D Results
Results of our introduced methodology are reported in Table II. Specifically, this table reports the results, i.e., Precision, Recall, and F1-score, per class and the overall Accuracy. We observe that our introduced model achieves F1-score greater than 0.99900 in terms of the Benign, TCP Connect Scan, ICPM flood, and UDP flood classes. Specifically, attacks of ICPM flood and UDP flood are perfectly recognized by our model achieving 1 across all metrics. F1-score greater than (or equal to) 0.99800 (but lower than 0.99900) is yielded in terms of SYN Scan, SYN flood, and HTTP flood. Finally, we observe that our model performs not so well in recognizing UDP Scan and Slow rate DoS reaching F1-score up to 0.99796 and 0.99735 respectively.
TABLE II: Results of our proposed approach.
@@
Table III reports results of our method in comparison with state-of-the-art approaches. As one can observe, our proposed model outperforms existing studies in Precision, Recall, and Accuracy by 0.00044-0.00875, 0.00043-0.01577, and 0.00148-0.00835 respectively. Our method outperforms, also, existing studies apart from Customized CNN (N=100) [23] in F1-score by 0.01228-0.00194. However, it must be noted that the study of [23] outperforms our study by a small margin of 0.00001. In the case of our study, we also report weighted F1-score, in order to take data imbalance into account. As one can observe, our proposed model reaches weighted F1-score up to 0.99958. Therefore, our study considering conditional-input components enjoys the benefits of lower training and inference time among else and achieves comparable performance to state-of-the-art results.
TABLE III: Performance comparison among proposed models and baselines.
@@ V-E Ablation Study
In this section, we perform a series of ablation experiments to explore the effectiveness of the proposed architecture. Results of the ablation study are reported in Table IV.
@@
Secondly, we remove the MoE layer and use a dense layer instead. To be more specific, the output vector of the CNN layers ($128D$) is passed through a dense layer consisting of 9 units (output layer). Results showed that Accuracy presented a decrease from 0.99958 to 0.99829. Specifically, F1-score of the Benign, SYN Scan, ICPM flood, SYN flood, HTTP flood, and Slow rate DoS presented a decrease of 0.00003, 0.00069, 0.00216, 0.00026, 0.00532, and 0.01052 respectively.
Finally, we remove both CNN and MoE layers, and train a deep neural network consisting of an input layer (78 units) and an output layer (9 units). Findings showed that Accuracy was decreased by 0.00802. F1-scores of all the classes, except for SYN Scan, presented a decrease. Specifically, F1-score of Benign, TCP Connect Scan, UDP Scan, ICPM flood, UDP flood, SYN flood, HTTP flood, and Slow rate DoS presented a decrease of 0.00515, 0.00082, 0.00063, 0.00324, 0.00001, 0.0123, 0.02683, and 0.0451 respectively.
TABLE IV: Ablation Study.
@@
Figure 3: Ablation Study. Varying the number of experts and top-k experts.

### v2

@@ Abstract
The advent of 6G/NextG networks comes along with a series of benefits, including extreme capacity, reliability, and efficiency. 6G/NextG networks must be equipped with advanced Artificial Intelligence algorithms, in order to evade new security threats. Existing studies on the intrusion detection task are based on deep neural networks consisting of static components, which are not conditional on the input. This limits their representation power and efficiency. To resolve these issues, we present the first study integrating Mixture of Experts (MoE) for identifying malicious traffic. Specifically, we use network traffic data and convert the 1D array of features into a 2D matrix. Next, we pass this matrix through convolutional neural network (CNN) layers followed by batch normalization and max pooling layers. After this, a sparsely gated MoE layer is used. This layer consists of a set of experts (dense layers) and a router, where the router assigns weights to the output of each expert. Sparsity is achieved by choosing the most relevant experts of the total ones. Finally, we perform a series of ablation experiments to prove the effectiveness of our proposed model. Experiments are conducted on the 5G-NIDD dataset, a network intrusion detection dataset generated from a real 5G test network, and the NANCY dataset, which includes cyberattacks on O-RAN 5G Testbed Dataset. Results show that our introduced approach reaches Accuracy up to 99.96% and 79.59% in terms of the 5G-NIDD and NANCY datasets respectively. Findings also show that our proposed model achieves multiple advantages over state-of-the-art approaches.
@@ I Introduction
Existing studies train shallow machine learning classifiers, obtaining in this manner suboptimal performance and poor generalization. Recently, existing studies have moved their interests in converting network traffic data or packet-level data into images and then training convolutional neural networks (CNNs), pretrained CNNs (AlexNet, VGG19, ResNet), RNNs (LSTMs, BiLSTMs), and after that employing fully-connected layers being responsible for classification. However, these approaches depend on dense layers being responsible for doing everything. Network parameters are fixed during training, while inference is performed in a static manner, demanding in this way computational resources and increasing both training and inference time. On the contrary, literature review suggests that models conditioned on the input [5] enjoy a series of benefits, including efficiency, representation power, adaptiveness, compatibility. This is attributable to the fact that these models selectively activate model components. Mixture of Experts belong to the category of the input-conditional computation models. Specifically, MoE were originally proposed in [6] and since then they have found applications in a series of domains [7, 8, 9].
To address the aforementioned limitations, we present the first study integrating MoE layers into a deep neural network in the task of intrusion detection in 5G networks. Specifically, we use network traffic data consisting of a feature set, which corresponds to a 1D array. We reshape this array into a matrix, which is then given as input to CNN, max-pooling, and batch normalization layers. After this, we use a sparsely-gated MoE layer [9], which applies different subsets of layers (experts) and activates only a selected subset of experts, i.e., the k most relevant ones, during each forward pass. Experiments are performed on two publicly available datasets, namely the 5G-NIDD dataset [10] and the O-RAN 5G Testbed NANCY dataset [11]. Results demonstrate that the proposed method achieves notable benefits over state-of-the-art approaches.
@@
  • We perform our experiments on two publicly available datasets related to 5G networks.
@@ III Datasets
Data of 5G-NIDD are extracted from two base stations. The authors in [10] have made available a csv file, which corresponds to the combined dataset, consisting of network flow data. 5G-NIDD dataset contains two attack categories, namely DoS/DDoS and Port Scan attacks. In terms of the DoS/DDoS attacks, the following types are included: ICMP flood, UDP flood, SYN flood, HTTP flood, and Slow rate DoS. Regarding the Port Scan attack, the following types are included: SYN Scan, TCP Connect Scan, and UDP Scan.
III-A1 Handling Missing Values
@@ III-A2 Designing the feature set
TABLE I: Description of features used (5G-NIDD dataset)
@@
III-B NANCY
NANCY dataset is a newly collected dataset. For constructing this dataset, a malicious user carries out cyberattacks against various services running in the main operator and the micro-operator. This dataset contains the following types of attacks, Reconnaissance Attack, TCP Scan, SYN Scan, SYN Flood, HTTP Flood, and Slow-rate DoS. The authors have used CICFlowMeter [^5] https://www.unb.ca/cic/research/applications.html software for obtaining the network flows. Due to data imbalance issues, we remove the Slowrate DoS and Reconnaissance Attack from our dataset. Specifically, the NANCY dataset includes 3,553 samples correponding to Slowrate DoS and 2,044 samples corresponding to Reconnaissance Attack.
Contrary to the 5G-NIDD dataset, the NANCY dataset does not include any missing values.
Designing the feature set
We keep only the numerical features. The complete feature set is reported in Table II. In total, we use 72 features. Min-Max Scaling is applied, ensuring values fall between 0 and 1. This 1D array is transformed into a 2D matrix, i.e., $6\times 12$, and is given as input to CNN layers, which are described in detail in Section IV.
TABLE II: Description of Features used (NANCY dataset)
Feature Description Feature Description
Flow Duration Duration of the flow in microseconds Tot Fwd Pkts Total number of packets in the forward direction
Tot Bwd Pkts Total number of packets in the backward direction TotLen Fwd Pkts Total length of packets in the forward direction
TotLen Bwd Pkts Total size of packets in the backward direction Fwd Pkt Len Max Maximum size of forward packets
Fwd Pkt Len Min Minimum size of forward packets Fwd Pkt Len Mean Mean size of forward packets
Fwd Pkt Len Std Standard deviation size of packet in forward direction Bwd Pkt Len Max Maximum size of packets in backward direction
Bwd Pkt Len Min Minimum size of backward packets Bwd Pkt Len Mean Mean size of backward packets
Bwd Pkt Len Std Standard deviation size of packet in backward direction Flow Byts/s Number of flow bytes per second
Flow Pkts/s Number of flow packets per second Flow IAT Mean Mean time between two packets sent in the flow
Flow IAT Std Standard deviation time between two packets sent in the flow Flow IAT Max Maximum time between two packets sent in the flow
Flow IAT Min Minimum time between two packets sent in the flow Fwd IAT Tot Total time between two packets sent in the forward direction
Fwd IAT Mean Mean time between two packets sent in the forward direction Fwd IAT Std Standard deviation time between two packets sent in the forward direction
Fwd IAT Max Maximum time between two packets sent in the forward direction Fwd IAT Min Minimum time between two packets sent in the forward direction
Bwd IAT Tot Total time between two packets sent in the backward direction Bwd IAT Mean Mean time between two packets sent in the backward direction
Bwd IAT Std Standard deviation time between two packets sent in the backward direction Bwd IAT Max Maximum time between two packets sent in the backward direction
Bwd IAT Min Minimum time between two packets sent in the backward direction Fwd Header Len Total header length in forward direction
Bwd Header Len Total header length in backward direction Fwd Pkts/s Number of forward packets per second
Bwd Pkts/s Number of backward packets per second Pkt Len Min Minimum packet length
Pkt Len Max Maximum packet length Pkt Len Mean Mean packet length
Pkt Len Std Standard deviation of packet lengths Pkt Len Var Variance of packet lengths
FIN Flag Cnt Number of packets with FIN flag SYN Flag Cnt Number of packets with SYN flag
RST Flag Cnt Number of packets with RST flag PSH Flag Cnt Number of packets with PSH flag
ACK Flag Cnt Number of packets with ACK flag URG Flag Cnt Number of packets with URG flag
CWE Flag Count Number of CWE flag packets ECE Flag Cnt Number of packets with ECE flag
Down/Up Ratio Ratio of bytes sent in backward and forward directions Pkt Size Avg Average packet size
Fwd Seg Size Avg Average segment size in forward direction Bwd Seg Size Avg Average segment size in backward direction
Fwd Byts/b Avg Average number of bytes per bulk in forward direction Fwd Pkts/b Avg Average number of packets per bulk in forward direction
Fwd Blk Rate Avg Average bulk rate in forward direction Bwd Byts/b Avg Average bytes per bulk in backward direction
Bwd Pkts/b Avg Average packets per bulk in backward direction Bwd Blk Rate Avg Average bulk rate in backward direction
Subflow Fwd Pkts Total packets in forward subflows Subflow Fwd Byts Total bytes in forward subflows
Subflow Bwd Pkts Total packets in backward subflows Subflow Bwd Byts Total bytes in backward subflows
Init Fwd Win Byts Initial window bytes in forward direction Init Bwd Win Byts Initial window bytes in backward direction
Fwd Act Data Pkts Number of forward active data packets Fwd Seg Size Min Minimum segment size in forward direction
Active Mean Mean active time Active Std Std of active time
Active Max Maximum activity time Active Min Minimum activity time
Idle Mean Mean idle time Idle Std Std of idle time
Idle Max Maximum idle time Idle Min Minimum idle time
IV Methodology
Figure 1: Proposed Methodology
(a) CNN Architecture ($f_{\theta}$)
(b) CNN cell structure
Figure 2: Description of $f_{\theta}$
IV-A Convolutional Neural Networks
@@
We design a convolutional neural network architecture to obtain a representation vector from input $X$. The CNN component of our methodology is illustrated in Fig. 2a.
@@
The structure of the CNN cell is illustrated in Fig. 2b. We place 4 CNN cell in a row with 16, 32, 64, and 128 filters respectively. Kernel size and padding are set equal to 3 and 1 respectively. In terms of the max pooling layer, we set both kernel size and stride equal to 2. Each CNN cell, except for the final one consists of an 1D Convolutional, Batch normalization, ReLU activation, and Max-Pooling layer. The final CNN cell does not include a max pooling layer.
@@
IV-B Sparse Mixture of Experts
@@ V-A Baselines
In terms of the 5G-NIDD dataset, we compare our method with the following research studies:
@@
Since NANCY dataset is a newly collected dataset, there are not research studies performing their experiments on this dataset.
@@ V-B Experimental Setup
We use PyTorch [34] for performing our experiments. All experiments are performed on a NVIDIA A100 80GB PCIe GPU.
5G-NIDD
We use 128 (value of $n$) experts. We keep the 32 (value of $k$) most relevant experts. We set $\alpha$ of Eq. 10 equal to 0.1. We use a batch size of 1024. We divide the dataset into a train and test set (60-40%) in a stratified manner. We train the models for a maximum of 40 epochs.
NANCY
We use 64 (value of $n$) experts. We keep the 32 (value of $k$) most relevant experts. We set $\alpha$ of Eq. 10 equal to 0.1. We use a batch size of 256. Also, we apply class weights. We divide the dataset into a train and test set (70-30%) in a stratified manner. We train the models for a maximum of 40 epochs.
@@ V-C Evaluation Metrics
Regarding 5G-NIDD, we also report weighted F1-score, since we have used an imbalanced dataset.
@@ V-D Results
Results on the 5G-NIDD dataset
Results of our introduced methodology are reported in Table III. Specifically, this table reports the results, i.e., Precision, Recall, and F1-score, per class and the overall Accuracy. We observe that our introduced model achieves F1-score greater than 0.99900 in terms of the Benign, TCP Connect Scan, ICMP flood, and UDP flood classes. Specifically, attacks of ICMP flood and UDP flood are perfectly recognized by our model achieving 1 across all metrics. F1-score greater than (or equal to) 0.99800 (but lower than 0.99900) is yielded in terms of SYN Scan, SYN flood, and HTTP flood. Finally, we observe that our model performs not so well in recognizing UDP Scan and Slow rate DoS reaching F1-score up to 0.99796 and 0.99735 respectively.
TABLE III: Results of our proposed approach (5G-NIDD)
@@
Table IV reports results of our method in comparison with state-of-the-art approaches. As one can observe, our proposed model outperforms existing studies in Precision, Recall, and Accuracy by 0.00044-0.00875, 0.00043-0.01577, and 0.00148-0.00835 respectively. Our method outperforms, also, existing studies apart from Customized CNN (N=100) [24] in F1-score by 0.01228-0.00194. However, it must be noted that the study of [24] outperforms our study by a small margin of 0.00001. In the case of our study, we also report weighted F1-score, in order to take data imbalance into account. As one can observe, our proposed model reaches weighted F1-score up to 0.99958. Therefore, our study considering conditional-input components enjoys the benefits of lower training and inference time among else and achieves comparable performance to state-of-the-art results.
TABLE IV: Performance comparison among proposed models and baselines on the 5G-NIDD dataset
@@
Results on the NANCY dataset
Table V presents the performance metrics of the proposed approach on the NANCY dataset, including Accuracy, Precision, Recall, and F1 Score for five traffic classes: Benign, SYN Scan, TCP Connect Scan, SYN flood, and HTTP flood. The overall accuracy of the model across all classes is 0.79592, indicating that approximately 79.6% of instances were correctly classified. Among the individual classes, the benign class achieved high performance, with a precision of 0.99637, recall of 0.97739, and F1 score of 0.98679, suggesting that normal traffic was accurately detected with minimal misclassifications.
In contrast, the model struggled to effectively detect SYN Scan traffic, achieving the lowest performance metrics with a precision of 0.59807, recall of 0.39572, and F1 score of 0.47629. This indicates a significant number of false negatives, pointing to challenges in identifying this type of scan-based intrusion. TCP Connect Scan performed moderately better, with a precision of 0.64581, recall of 0.74404, and an F1 score of 0.62969, reflecting a more balanced but still limited detection capability.
On the other hand, the approach showed excellent performance in detecting flooding attacks. SYN flood and HTTP flood traffic were classified with high precision and recall values—0.99865 and 0.98469 for SYN flood, and 0.99787 and 0.99760 for HTTP flood, respectively—resulting in F1 scores of 0.99162 and 0.99774. These results demonstrate the model’s robustness in identifying high-volume, aggressive attacks, while also highlighting the need for improvement in detecting subtler, low-rate intrusion attempts such as SYN scans.
TABLE V: Results of our proposed approach on the NANCY dataset
Class Accuracy Precision Recall F1 Score
Benign 0.79592 0.99637 0.97739 0.98679
SYN Scan 0.59807 0.39572 0.47629
TCP Connect Scan 0.64581 0.74404 0.62969
SYN flood 0.99865 0.98469 0.99162
HTTP flood 0.99787 0.99760 0.99774
@@ V-E Ablation Study
In this section, we perform a series of ablation experiments to explore the effectiveness of the proposed architecture. Results of the ablation study are reported in Table VI.
5G-NIDD
@@
Secondly, we remove the MoE layer and use a dense layer instead. To be more specific, the output vector of the CNN layers ($128D$) is passed through a dense layer consisting of 9 units (output layer). Results showed that Accuracy presented a decrease from 0.99958 to 0.99829. Specifically, F1-score of the Benign, SYN Scan, ICMP flood, SYN flood, HTTP flood, and Slow rate DoS presented a decrease of 0.00003, 0.00069, 0.00216, 0.00026, 0.00532, and 0.01052 respectively.
Finally, we remove both CNN and MoE layers, and train a deep neural network consisting of an input layer (78 units) and an output layer (9 units). Findings showed that Accuracy was decreased by 0.00802. F1-scores of all the classes, except for SYN Scan, presented a decrease. Specifically, F1-score of Benign, TCP Connect Scan, UDP Scan, ICMP flood, UDP flood, SYN flood, HTTP flood, and Slow rate DoS presented a decrease of 0.00515, 0.00082, 0.00063, 0.00324, 0.00001, 0.0123, 0.02683, and 0.0451 respectively.
TABLE VI: Ablation Study (5G-NIDD dataset)
@@
Figure 3: Ablation Study (5G-NIDD dataset). Varying the number of experts and top-k experts.
NANCY
Firstly, we set $L_{load}=L_{importance}=0$. Our proposed model (Table V) achieved an overall accuracy of 0.79592, whereas this model (Table VII) shows a slightly lower accuracy of 0.78614. While the difference in overall accuracy is minor, a closer examination of class-wise performance reveals more insightful distinctions. For the Benign class, the first model slightly outperforms the second, with an F1 score of 0.98679 compared to 0.98389, indicating marginally better consistency in identifying normal traffic. In the case of SYN Scan detection, the first model again shows superior performance across all metrics: precision (0.59807 vs. 0.57368), recall (0.39572 vs. 0.35541), and F1 score (0.47629 vs. 0.43890). Although both models struggle with detecting SYN scans, the first model demonstrates a relatively better capacity to identify this type of stealthy intrusion. For TCP Scan (Connect Scan), the second model performs almost identically to the first in terms of recall (0.74533 vs. 0.74404) but has slightly lower precision (0.52975 vs. 0.64581) and F1 score (0.61932 vs. 0.62969). This suggests that while both models detect similar quantities of true positives, the first model does so with fewer false positives, making it more reliable overall for this class. Regarding SYN flood detection, the F1 scores are nearly identical (0.99070 vs. 0.99162), indicating both approaches are highly effective for this attack type with negligible differences. Finally, for HTTP flood detection, one can observe a drop in F1-score from 0.99774 to 0.99718.
Next, we remove the MoE layer and observe differences in performance. Our proposed model demonstrates better overall accuracy (0.79592 vs. 0.77876). For benign traffic, both models perform similarly, but the proposed model achieves marginally higher precision and F1-score, indicating slightly more reliable identification of normal activity. In detecting SYN Scan attacks, the model without the MoE layer shows a significantly higher recall (0.83337 vs. 0.39572), meaning it captures more true positives, though model with MoE layer maintains better precision, resulting in a higher overall balance in performance. For TCP Scan, the introduced model clearly outperforms model without MoE with much higher recall (0.74404 vs. 0.23221) and F1 score (0.62969 vs. 0.33354), highlighting proposed model’s superior ability to detect this attack type. Both models perform exceptionally well in detecting SYN flood and HTTP flood attacks, with model without MoE layer having slightly better recall in SYN flood detection, though the differences are minimal. Overall, our proposed model incorporating both CNN and MoE provides a more balanced and consistent performance across all traffic classes.
Finally, we use a neural network, i.e., 72 units (input layer), 36 units (hidden layer), 5 units (output layer). Results show that CNN+MoE outperforms the dense-layer model in terms of overall accuracy (0.79592 vs. 0.76558), indicating better general classification performance. For benign traffic, CNN+MoE achieves slightly higher precision, recall, and F1 score, suggesting it identifies normal behavior more accurately. In detecting SYN Scan attacks, the dense-layer model demonstrates notably better recall (0.80135 vs. 0.39572), meaning it detects more true positives, though CNN+MoE maintains higher precision. For TCP Scan detection, both models perform weakly, but CNN+MoE again provides a better balance, with significantly higher recall (0.74404 vs. 0.23689) and F1 score (0.62969 vs. 0.33029). In flood-based attacks, both models perform very well; however, CNN+MoE slightly edges out the dense-layer model in HTTP flood detection, while both are comparably strong for SYN flood. Overall, CNN+MoE exhibits more consistent and balanced performance across all classes, particularly in scan-type attack detection.
TABLE VII: Ablation Study (NANCY dataset)
Evaluation metrics
Architecture Accuracy Precision Recall F1-score
Set importance and load loss to zero
Benign 0.78614 0.99486 0.97317 0.98389
SYN Scan 0.57368 0.35541 0.43890
TCP Scan 0.52975 0.74533 0.61932
SYN flood 0.99754 0.98395 0.99070
HTTP flood 0.99802 0.99635 0.99718
Remove MoE layer
Benign 0.77876 0.99355 0.97815 0.98579
SYN Scan 0.51127 0.83337 0.63374
TCP Scan 0.59182 0.23221 0.33354
SYN flood 0.99496 0.98662 0.99077
HTTP flood 0.99817 0.99589 0.99703
Simple input and dense layer
Benign 0.76558 0.97819 0.95929 0.96865
SYN Scan 0.49875 0.80135 0.61484
TCP Scan 0.54528 0.23689 0.33029
SYN flood 0.99721 0.97579 0.98638
HTTP flood 0.98796 0.98394 0.98595